Sally travels by car from one city to another. She drives for 27.0 min at 72.0 km/h, 54.0 min at 33.0 km/h, and 39.0 min at 74.0 km/h, and she spends 8.0 min eating lunch and buying gas.

(a) Determine the average speed for the trip.
1
Expert's answer
2012-03-27T10:00:14-0400
First, let's admit that

27.0 min = 27/60 h = 9/20 h;
54.0 min = 54/60 h = 9/10 h;
39.0 min = 39/60 h = 13/20 h.

Then, time spent on a trip is

T = 9/20 + 9/10 + 13/20 = 2h.

Total distance is

D = 9/20*72 + 9/10*33 + 13/20*74 = 551/5 km.

So, average speed for the trip is

V = D/T = 551/5/2 = 55.1 km/h.